### v1

@@ 3 Simulation Framework of CloudNativeSim
To ensure the simulator’s flexibility and broaden the scope of modeling for potential use cases, CloudNativeSim has been developed using a discrete-event simulation approach. As illustrated in Figure 2, the architecture of CloudNativeSim is structured in layers: (i) The User Layer is dedicated to user interaction, emphasizing a streamlined interface to significantly enhance the user experience. (ii) The Event Layer is responsible for scheduling discrete events, enabling the simulator to handle a broader spectrum of scenarios efficiently. (iii) The Entity Layer consists of various entities representing the modeling of microservice-based architecture within the simulator; this layer also serves as the foundation for executing discrete events. (iv) The bottom layer integrates core components from the CloudSim toolkit 8, providing a robust and efficient simulation environment.
@@ 3.1 User Layer: file registry and QoS feedback
In the rapidly evolving field of cloud-native applications, optimizing user interaction is particularly important 18. At the user level, our primary task is to accurately understand user needs and the specific configurations required by the simulator. To enable users to tailor CloudNativeSim scenarios to their requirements, we offer flexible customization options and design a series of examples to facilitate testing and provide instructive demonstrations.
The simulator provides extensible interfaces for essential policies, including horizontal and vertical scaling, allowing users to adjust these policies’ parameters based on specific service requirements and scenarios. This approach ensures that CloudNativeSim maintains a high level of customization and flexibility while enhancing the user experience.
@@
Furthermore, in order to lower the barrier for users to describe configurations of cloud-native applications, CloudNativeSim employs a file-based method for entity registration. Figure 3 illustrates a demo of the entity registry in CloudNativeSim, detailing configurations in both JSON and YAML formats. In the JSON section, APIs and services are defined. The API "POST /orders" has a weight of 1.0, while the "orders" service is labeled "orders" and calls services like "orders-db" and "carts". The YAML configuration specifies instance settings, including the "orders" pod, its labels, replicas, size, and bandwidth settings. It also details resource requests and limits for shares and RAM. These configurations provide a clear framework for managing APIs, services, and instances, ensuring efficient resource allocation and robust simulation of cloud-native environments. By focusing solely on the content of configuration files, users can operate the simulator without requiring an in-depth understanding of its mechanisms 19.
To more effectively manage and analyze the performance of cloud-native applications, CloudNativeSim offers extensive feedback on QoS. This includes traditional instance-based metrics such as response time and utilization, as well as more detailed service-oriented metrics like Service Level Objective (SLO) violation rate and request-based metrics such as Queries Per Second (QPS) and detailed response time. These metrics are crucial for performance evaluation and ensuring adherence to predefined service standards.
@@ 3.2 Event Layer: the core design for scheduling
  • Derivative: When a service completes the cloudlets assigned to it, it automatically derives new cloudlets to next nodes along the service chain. This process ensure the number of cloudlets within the simulator to vary with the number of requests and services, and it also forms the basis for calculating response times using the critical path.
@@ 4 Modeling of Cloud-native Applications
To comprehensively simulate the performance of a cloud-native application, it is necessary to model its internal structure. During the modeling process, the simulator must address the challenges posed by the heterogeneity and high dynamism inherent in microservice-based architecture. Heterogeneity refers to services operate in different code environments and are mapped to various instances, forming a complex entity communication network. Additionally, high dynamism, driven by varying request workloads, complicates the prediction of resource and state changes.
In response to these challenges, CloudNativeSim employs a novel modeling approach, as illustrated in Figure 5. This approach consists of three key components: (i) The service graph modeling constructs dependencies between services. It effectively describes the communication topology among heterogeneous entities. (ii) The cloudlet scheduler provides multi-level queues to manage cloudlets. This ensures flexible resource and status updates within the system. (iii) The client request processing module simulates dynamic request generation and calculates response times by updating the critical path. This captures the dynamic interactions between cloud-native applications and clients.
@@ 4.1.1 Service Dependency Modeling based on DAG
Given that communications between services is unidirectional and lacks cyclic calls, using a DAG structure is an efficient approach for modeling service dependencies 20, 21, 22. This structure enables more accurate and detailed simulations of service interactions and dependencies, providing valuable insights into system performance and potential failure points. By utilizing DAG, CloudNativeSim can help developers and researchers better understand the complex dynamics of cloud-native applications, which leads to optimized design and enhanced strategies.
@@
Figure 6 presents a simple DAG to depict the service dependencies: Service A calls Services B and C, which in turn call Service D. Each service comprises many code segments represented by rectangles of different colors. The DAG structure includes multiple service chains corresponding to APIs, and client requests will propagate along these chains, forming the inter-service communications and eventually responding to the clients. These chains can be recursively defined according to different source services. Figure 7 shows the built-in data structures of CloudNativeSim, representing a bidirectional service hierarchy. By retrieving these tables, our simulator can quickly derive tasks between services in the graph and update metrics in reverse.
Figure 7: The data strucure of bidirectional service hierarchy table in CloudNativeSim. The left table shows the forward lookup of service dependencies where each service points to its dependent services. The right table illustrates the reverse lookup where each service lists the services that depend on it.
@@
By modeling the service graph through a DAG structure from both horizontal and vertical perspectives, CloudNativeSim offers a clear and detailed representation of service dependencies. This approach enhances user insights into cloud-native applications, making internal changes more transparent and easier to understand.
@@ 4.1.2 Inter-service Communication Modeling based on RpcCloudlet
An efficient inter-service communication mechanism can significantly reduce request response times 23. The widely accepted gRPC mechanism  [^1] https://grpc.io/ enhances this efficiency 24. To achieve similar results, CloudNativeSim introduces RpcCloudlet to model the communication process between services. RpcCloudlet is a extension of Cloudlet, which is the most basic task unit designed in CloudSim.
Although the role of RpcCloudlet in CloudNativeSim is similar to that of procedure calls, there are distinct differences between the physical system and the simulator. Figure 8(a) shows the Jaeger [^2] https://www.jaegertracing.io/ monitoring results of the sockshop application. When a request to the "Get /login" API arrives, it triggers a series of endpoints (represented by ovals). These endpoints are associated with different services, forming multiple service chains. In contrast, Figure 8(b) illustrates a different invocation scenario in the simulator: requests are not directly distributed to service endpoints. Instead, multiple cloudlets are created by services to represent the communications. Furthermore, the task lengths in the actual system are typically unknown and independent. To simulate this characteristic, we introduce a Gaussian distribution to model their lengths when creating cloudlets.
@@
Additionally, for large-scale applications, the heterogeneity of code logic makes it challenging for users to determine how many tasks a single request will trigger within the current service. Additionally, requests may generate different types of tasks, resulting in varying task sizes and, consequently, differing cloudlet lengths. In such scenarios, it is impractical to require users to input precise cloudlet configurations at the start, as traditional simulators do. Instead, the simulator should automatically configure the number and size of cloudlets.
To meet this requirement, the number of cloudlets will automatically increase as inter-service communications grow. After the current service processes its cloudlets, it will check for the existence of subsequent services in the current chain and attempt to create and derive new cloudlets for them. This multi-cloudlet model retains the advantages of RPC, as the failure of a single cloudlet does not block the execution of other tasks, thereby enhancing the system’s reliability and dynamism.
In summary, the inter-service communication modeling approach aforementioned offers several advantages: (i) It simplifies user operations, eliminating the need for manual configuration of each cloudlet in detail and enabling dynamic modeling with minimal parameters. (ii) It offers strong extensibility. The random distribution introduced during cloudlet creation provides significant flexibility, accommodating a wider range of application scenarios. (iii) It introduces common task units to address the issue of code heterogeneity in microservice-based architectures. (iv) It is benefit to maintain the dynamism of inter-entity communication.
@@ 4.3 Processing of Client Requests
In the simulation of cloud-native applications, user requests serve as the core entry point, driving all scheduling activities. CloudNativeSim treats each service within a microservice-based architecture as a set of functions, acting as the top-level unit for handling user requests. Consequently, processing these requests becomes the primary objective of the simulation. Compared to traditional solutions, this request-oriented simulation emphasizes updating and providing feedback on request metrics. To achieve this, we need to describe two critical components in CloudNativeSim: request generation and response time calculation.
4.3.1 Request Generation Modeling based on Queuing Theory
To dynamically simulate the arrival of requests, CloudNativeSim employs a more versatile approach in which the arrival rate is dynamically dictated by the system’s parameters. The Request Generator component simulates user request arrivals based on specified parameters, such as the number of clients, growth rate, and waiting time intervals.
@@ 4.3.1 Request Generation Modeling based on Queuing Theory
Input: $N_{c}$ (number of clients), $v$ (velocity factor), $p$ (pause time range), $\{w\}$ (weight set), $apiList$ (list of APIs), $timeLimit$ (maximum time) or $numLimit$ (maximum number of requests)
@@
When designing the Request Generator component, we are inspired by Locust  [^3] https://locust.io/ and aimed to simplify it for simulation purposes. Algorithm 4.3.1 describes the logic of this component: at the beginning of the simulation, users need to specify several parameters: the final number of clients $N_{c}$, client growth rate $v$, request weight for each interface $w$ (default is 1), user request waiting span $p$, time limit $timeLimit$, or quantity limit $numLimit$. The request generation process is divided into the following steps:
@@
4.3.2 Response Time Modeling based on Critical Path
To accurately simulate the response time of requests within the system, the modeling based on critical path is efficient as it identifies the longest sequence of dependent tasks, determining the maximum completion time for a request 21. This approach helps users find out bottlenecks and optimize resource allocation, thereby enhancing the practicality and efficiency of CloudNativeSim.
Algorithm 2 Response Time Updating Algorithm based on Critical Path
@@ 4.3.2 Response Time Modeling based on Critical Path
Algorithm 4.3.2 illustrates the response time updating process based on the critical path modeling in CloudNativeSim. Initially, the algorithm sets the response time to zero and determines the service chain based on the API specified in the request (lines 3-4). For each path in the service chain, the algorithm initializes the current delay to zero and then iterates through each node in the path, accumulating the delays (lines 5-8). To describe this process mathematically, let $P$ denote the set of all paths in the service chain, and $N_{p}$ denote the set of nodes in path $p\in P$. The current delay for each path $p$, denoted as $D_{p}$, is computed by summing the delays of all nodes in the path:
@@
where $\text{delay}(n)$ represents the delay associated with node $n$.
If the delay of the current path exceeds the previously recorded maximum response time, the algorithm updates the response time and records the current path as the critical path (lines 10-12). After examining all paths, the algorithm updates the service graph with the identified critical path and its corresponding response time (lines 14-15). The critical path $CP$ is then updated to the path $p$ that has the maximum delay:
@@
Based on this modeling approach, CloudNativeSim can efficiently simulate service request transmissions while providing real-time updates and optimization suggestions for the critical path. It ensures that the response time of requests is accurately calculated based on the longest service execution path, which is crucial for identifying performance bottlenecks and optimizing system response times.

### v2

@@ 3 Simulation Framework of CloudNativeSim
To ensure the simulator’s flexibility and broaden the scope of modeling for potential use cases, CloudNativeSim has been developed using a discrete-event simulation approach. As illustrated in Figure 2, the architecture of CloudNativeSim is structured in layers: (i) The User Layer is dedicated to user interaction, emphasizing a streamlined interface to significantly enhance the user experience. (ii) The Event Layer is responsible for scheduling discrete events, enabling the simulator to handle a broader spectrum of scenarios efficiently. (iii) The Entity Layer consists of various entities representing the modeling of microservice-based architecture within the simulator; this layer also serves as the foundation for executing discrete events. (iv) The bottom layer integrates core components from the CloudSim toolkit 9, providing a robust and efficient simulation environment.
@@ 3.1 User Layer: file registry and QoS feedback
In the rapidly evolving field of cloud-native applications, optimizing user interaction is particularly important 19. At the user level, our primary task is to accurately understand user needs and the specific configurations required by the simulator. To enable users to tailor CloudNativeSim scenarios to their requirements, we offer flexible customization options and design a series of examples to facilitate testing and provide instructive demonstrations. The user layer provides options including deployment specifications, logging preferences, policy configurations and cloudlet parameters. Users can also tailor the reporting format to output specific metrics to either terminal displays or CSV files for subsequent analysis.
@@
Furthermore, in order to lower the barrier for users to describe configurations of cloud-native applications, CloudNativeSim employs a file-based method for entity registration. Figure 3 illustrates a demo of the entity registry in CloudNativeSim, detailing configurations in both JSON and YAML formats. In the JSON section, APIs and services are defined. The API "POST /orders" has a weight of 1.0, while the "orders" service is labeled "orders" and calls services like "orders-db" and "carts". The YAML configuration specifies instance settings, including the "orders" pod, its labels, replicas, size, and bandwidth settings. It also details resource requests and limits for shares and RAM. These configurations provide a clear framework for managing APIs, services, and instances, ensuring efficient resource allocation and robust simulation of cloud-native environments. By focusing solely on the content of configuration files, users can operate the simulator without requiring an in-depth understanding of its mechanisms 20.
To more effectively manage and analyze the performance of cloud-native applications, CloudNativeSim provides dynamic QoS feedback through continuous monitoring and recording of performance metrics throughout the simulation runtime. This includes traditional instance-based metrics such as resource utilization, as well as detailed service-oriented metrics including interface-specific response times, Requests Per Second (RPS), and Service Level Objective (SLO) violation rates at both instance and system levels. All these metrics are displayed in real-time through terminal output, with configurable sampling frequencies to control reporting granularity. Users can further customize the reporting format by selecting specific metrics for display or exporting comprehensive logs to CSV files for detailed post-simulation analysis. These comprehensive monitoring capabilities enable thorough evaluation of cloud-native application behavior under various operational conditions, ensuring adherence to predefined service standards.
@@ 3.2 Event Layer: the core design for scheduling
  • Derivative: When a service completes the cloudlets assigned to it, it automatically derives new cloudlets to next nodes along the service chain. This process ensure the number of cloudlets within the simulator to vary with the number of requests and services, and it also forms the basis for calculating response time using the critical path.
@@ 4 Modeling of Cloud-native Applications
To comprehensively simulate the performance of a cloud-native application, it is necessary to model its internal structure. During the modeling process, the simulator must address the challenges posed by the heterogeneity and high dynamism inherent in microservice-based architecture. Heterogeneity refers to services operating in different code environments and are mapped to various instances, forming a complex entity communication network. Additionally, high dynamism, driven by varying request workloads, complicates the prediction of resource and state changes.
Terms Definitions Terms Definitions
$service$ A service is an independent unit that performs functionality and communicates with other services via APIs. $instance$ An instance refers to a specific performer of a service, such as a pod or container.
$task$ A task corresponds to a request that invokes multiple services and is completed by several cloudlets working together. The task length is the total computational workload measured in millions of instructions (MI). $cloudlet$ A cloudlet is a fundamental computational unit used for task processing in the simulation. The cloudlet length is also measured in MI.
$instanceList$ A collection of all active service instances in the system at a given time. $scalingList$ A dynamic list of instances marked for scaling operations (scale-out or scale-in).
$apiList$ A list of all available APIs exposed by services in the system. $\{w\}$ A set of weights representing the probability distribution for API selection by clients.
$P$ The set of all possible execution paths through the service chain, where each path represents a sequence of service invocations. $CP$ The critical path represents the longest execution sequence in terms of response time among all possible paths $P$.
$c_{i}$ The $i$-th client entity in the system that generates service requests according to defined patterns. $N_{c}$ The current total number of active clients in the system at time $t$.
$v$ The client growth rate, defining how $N_{c}$ changes over time ($v>0$ for growth). $p$ The inter-request interval range $p=[p_{0},p_{1}]$, where $p_{0}$ and $p_{1}$ represent the minimum and maximum waiting time between consecutive requests.
$timeLimit$ The maximum time period during which clients are allowed to send requests. $numLimit$ The maximum number of requests to be processed in the simulation.
$\lambda$ The request arrival rate (QPS), measuring the number of incoming requests per second. $R$ The cumulative total number of requests since the start of simulation until the current time point.
Table 2: Terms and definitions for cloud-native application modeling in CloudNativeSim.
To facilitate clear communication and understanding, we first introduce the key terms and their definitions used throughout this paper in Table 2. These terms form the foundation of our modeling approach and will be referenced extensively in subsequent sections.
In response to the above challenges, CloudNativeSim employs a novel modeling approach, which consists of three key components: (i) The service graph modeling constructs dependencies between services, effectively describing the communication topology among heterogeneous entities. (ii) The cloudlet scheduler provides multi-level queues to manage cloudlets, ensuring flexible resource and status updates within the system. (iii) The client request processing module simulates dynamic request generation and calculates response time by updating the critical path, capturing the dynamic interactions between cloud-native applications and clients.
4.1 Modeling of Service Graph
@@ 4.1.1 Service Dependency Modeling based on DAG
Given that communications between service communication is unidirectional and lacks cyclic calls, using a DAG structure is an efficient approach for modeling service dependencies 21, 22, 23. This structure enables more accurate and detailed simulations of service interactions and dependencies, providing valuable insights into system performance and potential failure points. By utilizing DAG, CloudNativeSim can help developers and researchers better understand the complex dynamics of cloud-native applications, which leads to optimized design and enhanced strategies.
@@
Figure 6 presents a simple DAG to depict the service dependencies: Service A calls Services B and C, which in turn call Service D. Each service comprises many code segments represented by rectangles of different colors. The DAG structure includes multiple service chains corresponding to APIs, and client requests will propagate along these chains, forming the inter-service communications and eventually responding to the clients. These chains can be recursively defined according to different source services. Figure 7 shows the built-in data structures of CloudNativeSim, representing service dependencies through a DAG implementation. By utilizing these reverse-linked lists, our simulator can efficiently track service interactions and trace call chains in both forward and reverse directions.
Figure 7: The data structure of service dependency tracking in CloudNativeSim. The left table implements forward lookup where each service maps to its dependent services. The right table implements reverse lookup where each service maps to services that call it.
@@
By modeling the service graph through a DAG structure, CloudNativeSim offers a clear and detailed representation of service dependencies. This approach enhances user insights into cloud-native applications, making internal changes more transparent and easier to understand.
@@ 4.1.2 Inter-service Communication Modeling based on RpcCloudlet
Inter-service communication efficiency serves as a critical determinant in modern distributed systems’ performance 24. The emergence of gRPC  [^2] https://grpc.io/ has revolutionized this domain by introducing a high-performance, language-agnostic communication protocol 25. Drawing inspiration from these advances, CloudNativeSim introduces RpcCloudlet as a novel modeling construct to simulate inter-service communications. This extension builds upon CloudSim’s fundamental Cloudlet abstraction while incorporating contemporary distributed computing paradigms.
The architectural distinction between physical systems and their simulated counterparts becomes evident when examining their respective communication patterns. Figure 8(a) illustrates this through Jaeger’s  [^3] https://www.jaegertracing.io/ distributed tracing visualization of the sockshop microservices application. In production environments, a "Get /login" API request initiates a cascade of service invocations, where each endpoint (represented as ovals) belongs to distinct microservices, forming intricate service dependency chains. Conversely, Figure 8(b) demonstrates the simulator’s approach, where inter-service communications are abstracted through dynamically generated cloudlets, offering a more granular representation of these complex interactions.
Our implementation marks a significant difference from CloudSim’s traditional static cloudlet allocation model. Where CloudSim employs upfront cloudlet definition and assignment to virtual machines—an approach that fails to capture the dynamic nature of modern microservices—our system introduces an innovative "dynamic derivation" mechanism. This approach generates cloudlets on-demand, precisely when their corresponding services are invoked, establishing a more realistic representation of service interactions. A distinguishing feature of this design is its ability to accommodate multiple cloudlet instances for a single service throughout the request lifecycle. As illustrated in Figure 8(b), the user-db service exemplifies this capability, spawning multiple cloudlet instances in response to various service invocations. This dynamic instantiation pattern closely aligns with the principles of gRPC communication, hence the designation "RPCCloudlet."
@@
To address the inherent complexity of request processing, our system implements a novel cloudlet management mechanism. Through empirical observation, we discovered that computational requirements exhibit statistical independence, making them amenable to modeling via Gaussian distribution. The cloudlet configuration framework incorporates user-definable parameters, including expected mean length, variance, and service-instance mapping specifications. Upon parameter initialization, the system autonomously derives new cloudlets as service invocations occur.
This comprehensive approach to inter-service communication modeling offers several compelling advantages: First, the cloudlet derivation process inherently reflects service dependencies, enhancing simulation transparency and interpretability. Second, the integration with a discrete event framework enables dynamic performance metric evolution, particularly in latency measurements, contributing to enhanced system adaptability. Third, the design preserves fundamental RPC characteristics, ensuring that individual cloudlet failures do not compromise overall system execution, thereby maintaining robust operational continuity and system resilience. These features collectively establish a more realistic and sophisticated simulation environment for modern distributed systems.
@@ 4.3 Processing of Client Requests
In the simulation of cloud-native applications, user requests serve as the core entry point, driving all scheduling activities. Consequently, processing these requests becomes the primary objective of the simulation. Compared to traditional solutions, this request-oriented simulation emphasizes updating and providing feedback on request-related metrics (including latency, RPS, and SLO violation rates). To achieve this, we need to describe two critical components in CloudNativeSim: request generation and response time calculation.
4.3.1 Request Generation Modeling
To dynamically simulate the arrival of requests, CloudNativeSim employs a more versatile approach in which the arrival rate is dynamically dictated by the system’s parameters. The Request Generator component simulates user request arrivals based on specified parameters, such as the number of clients ($N_{c}$), client growth rate ($v$), waiting time intervals ($p$), request weight for each API ($\{w\}$), list of APIs ($apiList$), maximum generation time ($timeLimit$), and maximum number of requests ($numLimit$). The client growth rate ($v$) represents the rate at which the number of clients increases over time, acting as a velocity factor that dictates how quickly new clients are added to the system. Waiting time intervals ($p$) define the range of time intervals that clients wait between sending consecutive requests, represented as an interval $p=[p_{0},p_{1}]$, where $p_{0}$ is the minimum wait time and $p_{1}$ is the maximum wait time. Each client will randomly select a wait time from this interval after sending a request. Request weight for each API ($\{w\}$) is a set of weights that represent the probability distribution for selecting different APIs when generating requests. Each weight corresponds to an API in the $apiList$, determining how likely it is for a particular API to be chosen when a client sends a request.
@@ 4.3.1 Request Generation Modeling
Input: $N_{c}$ (number of clients), $v$ (client growth rate), $p$ (waiting time intervals), $\{w\}$ (weight set), $apiList$ (list of APIs), $timeLimit$ (maximum generation time) and $numLimit$ (maximum number of requests)
@@
When designing the Request Generator component, we were inspired by Locust  [^4] https://locust.io/ and aimed to simplify it for simulation purposes. Algorithm 4.3.1 describes the logic of this component: at the beginning of the simulation, users need to specify several parameters: the final number of clients $N_{c}$, client growth rate $v$, request weight for each interface $w$ (default is 1), user request waiting span $p$, time limit $timeLimit$, or quantity limit $numLimit$. The request generation process is divided into the following steps:
@@
4.3.2 Response Time Calculation and Critical Path Identification
In cloud-native applications, the response time of a request is determined by the completion time of all its dependent tasks. CloudNativeSim employs a discrete event framework to track the execution of these tasks and calculate the response time. After obtaining the response time, the system identifies the critical path by backtracking through the service invocation chain, which helps users understand the performance bottlenecks and optimize resource allocation.
Figure 9: A demonstration for response time calculation and critical path identification. The red path (A→B→E→F) shows the identified critical path in service graph.
Figure 9 illustrates this process with a service invocation graph. The nodes (A-F) represent individual cloudlets, with edges showing their invocation relationships. The discrete event framework tracks the execution of each cloudlet, recording their start and completion time. The response time is determined when the last cloudlet (F in this case) completes its execution. Subsequently, given the one-to-one correspondence between cloudlets and services, the system performs a post-hoc analysis by backtracking from F through the cloudlets with the latest completion time, thereby identifying the critical path (highlighted in red as A→B→E→F) in the service graph. This critical path represents the sequence of services whose execution time directly impact the overall response time, which is particularly valuable for users in analyzing bottlenecks in cloud-native applications.
Algorithm 2 Response Time Calculation and Critical Path Identification
@@ 4.3.2 Response Time Calculation and Critical Path Identification
Algorithm 4.3.2 demonstrates the critical path identification process. The algorithm examines all possible paths in the service chain to identify the one with the maximum accumulated delay. Let $P$ denote the set of all paths in the service chain, and $N_{p}$ denote the set of nodes in path $p\in P$. For each path, the total delay $D_{p}$ is calculated as:
@@
where $\text{delay}(n)$ represents the actual execution time of node $n$ obtained from the discrete event simulation.
The critical path $CP$ is then determined as the path with the maximum total delay:
@@
This approach enables CloudNativeSim to provide valuable insights into service performance. By analyzing the critical path, users can identify which service invocations contribute most significantly to the overall response time. This information is particularly valuable for system optimization, as it highlights potential bottlenecks and suggests where performance improvements would be most effective.